# Classical spin liquids from frustrated Ising models in hyperbolic space

### v1

@@ Abstract
The behavior of many-body systems is crucially influenced by symmetry and geometry of the underlying real-space lattice, with band degeneracies, band topology, as well as frustration and related interference effects being key examples. Interest in recent years has extended to lattices embedded in curved space, hyperbolic space in particular, for a number of reasons: First, many condensed-matter paradigms established for Euclidean flat space are substantially modified in hyperbolic space, for instance band theory [1, 2, 3] and critical behavior [4, 5, 6]. Second, the holographic principle as realized by the AdS/CFT correspondence postulates a connection between interacting quantum field theories and gravity theories in higher dimensions [8, 7, 9], and lattice-regularized gravity theories lead to statistical-mechanics models on hyperbolic lattices [10, 11, 12, 13, 14]. Third, hyperbolic systems have been discussed in quantum information theory as candidates to efficiently encode qubits using toric code [15].
@@
Discretizations of curved space lead to the concept of tesselations: A uniform tessellation $\{p,q\}$ denotes a two-dimensional space-filling lattice of regular $p$-sided polygons where $q$ polygons meet at each lattice vertex. Such tessellations can naturally be embedded on two-dimensional surfaces of constant curvature $\rho$ [3, 35]. Depending on the sign of the characteristic $\chi=1/2-1/p-1/q$ we can distinguish between spherical ($\chi<0$), flat ($\chi=0$) and hyperbolic ($\chi>0$) tessellations which are embedded on the eponymous space. Geometric frustration requires $p$ to be odd. To generate maximal frustration, we focus on hyperbolic tessellations that are composed of triangles, $p=3$. Since the simplest realization of $\chi>0$ is given by $q=7$, we focus on the $\{3,7\}$ tessellation for most of this work.
Figure 1: Ising model on the $\{3,7\}$ lattice with type-A boundary and $N=85$ spins. (a) Snapshot spin configuration at low $T$, with black (white) dots indicating up (down) spins and red (blue) bonds indicating frustrated (unfrustrated) bonds. (b) Color-coded NN bond correlations $\langle\sigma_{i}\sigma_{j}\rangle$ at $T\!=\!0.2$, showing alignment along the boundary.
Uniform hyperbolic tessellations represent infinite lattices; they can be understood as limits of finite hyperbolic uniform tilings. These finite lattices have unique features that distinguish them from their planar counterparts: The number of sites grows exponentially with linear system size, leading to strong boundary contributions. Implementing periodic boundary conditions is not straightforward: Their possible unit cells are restricted by the Gauss-Bonnet theorem and involve the solution of lattice-specific diophantine equations [3, 6]; we therefore choose to work with open boundary conditions. There are a multitude of ways to construct finite lattices, themselves leading to different shapes of the lattice boundary. We choose the iterative algorithm of Refs. 12, 35, described in more detail in the supplemental material (SM) [36]. A sample lattice is shown in Fig. 1 as a planar graph using Poincaré projection; we refer to this type of boundary as type-A. Other boundary shapes will be discussed below.
Figure 2: Energy per site, $E/N$, as function of temperature $T$ for $\{3,7\}$ type-A systems of different size $N$, obtained from MC simulations. The inset shows the low-temperature part of the data.
@@
Sample results for internal energy, entropy, and specific heat of type-A systems are shown in Figs. 2 and 3, respectively; a snapshot low-$T$ configuration is in Fig. 1(a). The energy $E(T)/N$ reaches a constant value upon cooling below $T/J\lesssim 0.1$; this constant approaches $(-J)$ with increasing system size and is in agreement with the analytically determined ground-state energy as discussed below. The specific heat $C_{V}(T)/N$ displays a broad maximum around $T/J\approx 0.6$ whose height and shape is essentially independent of system size. At low temperatures we get enlarged statistical errors because of spin freezing.
@@
$S_{\rm res}=\ln(2)+\frac{1}{2}\operatorname{logdet}(A^{\prime}),$ (4)
which we have used to compute $S_{\rm res}$ for system sizes up to $N=6000$. The corresponding numerical results for $\{3,7\}$ type-A systems are shown in Fig. 6; here we have also included data for systems with partially completed layers. MC data for $s_{\rm res}$ are shown for comparison; these are consistent with the FKT data within error bars for cases where the same system size is covered. We note that the $N=232,617$ systems feature an odd number of plaquettes such that the FKT algorithm is not applicable [41]. Despite significant finite-size effects, the residual entropy clearly converges to a finite value for $N\to\infty$, which we estimate as $s_{\rm res}\equiv S_{\rm res}/N=0.102(2)$.
@@
where the number of vertices ($N$), edges ($N_{E}$) and plaquettes ($N_{P}$) of the lattice are related by Euler’s characteristic for planar graphs, $N-N_{E}+N_{P}=1$; this differs by unity from the literature since our definition of plaquettes excludes the outside. For our hyperbolic lattice we can obtain closed expressions for $N$, $N_{E}$, and $N_{P}$ based on its iterative construction. This eventually leads to the result [36]
@@
valid for large $n$ where $n$ denotes the number of lattice layers, and $\delta_{n}$ is one if $n$ is divisible by three and zero otherwise. The ground-state energy hence approaches $(-JN)$ for $N\to\infty$, in agreement with Fig. 2.
@@
We have also studied hyperbolic $\{3,7\}$ tilings with different boundary geometry. As detailed in the SM, a type-B lattice is obtained by adding one site, i.e. one outer triangle, to each boundary bond of a type-A lattice. This construction generates a family of type-B tilings. Our numerical data for type-B system indicate that they settle into an ordered ferrimagnetic ground state, with ferromagnetic order in each layer and antiferromagnetic alignment from layer to layer. The reason is energetic: Type-B systems have more boundary bonds than type-A systems, with a zigzag boundary shape, such that perfect antiferromagnetic order along the boundary coincides with ferromagnetic intralayer order, thus tipping the balance towards ordered states. The thermodynamic behavior of type-B systems is analyzed in detail in the SM [36].
@@
Among the directions for future research are the investigation of quantum effects towards triangular hyperbolic quantum spin liquids – this suggests for instance to consider quantum dimer models on the present lattices – as well as the study of classical spin liquids in hyperbolic vector spin models. More broadly, open questions concern how lattice frustration manifests in different curved-space geometries, e.g., with black holes included [13, 43], and whether the spin liquids found here have implications for gravitational theories.

### v2

@@ Abstract
The behavior of many-body systems is crucially influenced by symmetry and geometry of the underlying real-space lattice, with band degeneracies, band topology, as well as frustration and related interference effects being key examples. Interest in recent years has extended to lattices embedded in curved space, hyperbolic space in particular, for a number of reasons: First, many condensed-matter paradigms established for Euclidean flat space are substantially modified in hyperbolic space, for instance band theory [1, 2, 3] and critical behavior [4, 5, 6]. Second, the holographic principle as realized by the AdS/CFT correspondence postulates a connection between interacting quantum field theories and gravity theories in higher dimensions [8, 7, 9]. As a generalization, it has also been postulated that gravity theories discretized on hyperbolic tilings of two-dimensional hyperbolic space are dual to spin chains at the boundary [10, 11, 12, 13, 14], a program known as ‘discrete holography’. Third, hyperbolic systems have been discussed in quantum information theory as candidates to efficiently encode qubits using toric code [15].
@@
Discretizations of curved space lead to the concept of tesselations: A uniform tessellation $\{p,q\}$ denotes a two-dimensional space-filling lattice of regular $p$-sided polygons where $q$ polygons meet at each lattice vertex. Such tessellations can naturally be embedded on two-dimensional surfaces of constant curvature $\rho$ [3, 35]. The sign of the characteristic $\chi=1/2-1/p-1/q$ distinguishes between spherical ($\chi<0$), flat ($\chi=0$) and hyperbolic ($\chi>0$) tessellations, embedded on the eponymous space. Geometric frustration requires $p$ to be odd. To generate maximal frustration, we focus on hyperbolic tessellations that are composed of triangles, $p=3$. Since the simplest realization of $\chi>0$ is given by $q=7$, we focus on the $\{3,7\}$ tessellation for most of this work.
Figure 1: Snapshot low-temperature configurations of antiferromagnetic Ising models on $\{3,7\}$ lattices. Red/blue bonds indicate parallel/antiparallel spins on the respective bond. (a) System with type-A boundary with $N=29$ spins, showing a spin-liquid configuration with antiferromagnetic alignment along the boundary. (b) Type-B boundary with $N=50$ spins, showing an ordered ferrimagnet.
Uniform hyperbolic tessellations represent infinite lattices; they can be understood as limits of finite hyperbolic uniform tilings. These finite lattices have unique features that distinguish them from their planar counterparts: The number of sites grows exponentially with linear system size, leading to strong boundary contributions. Implementing periodic boundary conditions is not straightforward: Their possible unit cells are restricted by the Gauss-Bonnet theorem and involve the solution of lattice-specific diophantine equations [3, 6]; we therefore choose to work with open boundary conditions. There are a multitude of ways to construct finite lattices, themselves leading to different shapes of the lattice boundary. We choose the iterative algorithm of Refs. 12, 35, described in more detail in the supplemental material (SM) [36]. A sample lattice is shown in Fig. 1(a) as a planar graph using Poincaré projection; we refer to this type of boundary as type-A. Other boundary shapes will be discussed below.
@@
Results for entropy and specific heat of type-A systems are shown in Fig. 2; a snapshot low-$T$ configuration is in Fig. 1(a). The specific heat $C_{V}(T)/N$ displays a broad maximum around $T/J\approx 0.6$ whose height and shape is essentially independent of system size. At low temperatures we get enlarged statistical errors because of spin freezing. The energy $E(T)/N$, shown in the SM [36], reaches a constant upon cooling below $T/J\lesssim 0.1$; this constant approaches (J)(-J) with increasing system size, in agreement with the analytical result obtained below.
@@
$S_{\rm res}=\ln(2)+\frac{1}{2}\ln(\det A^{\prime}),$ (4)
which we have used to compute $S_{\rm res}$ for system sizes up to $N=6000$. The corresponding numerical results for $\{3,7\}$ type-A systems are shown in Fig. 5; here we have also included data for systems with partially completed layers. MC data for $s_{\rm res}$ are shown for comparison; these are consistent with the FKT data within error bars for cases where the same system size is covered. We note that the $N=232,617$ systems feature an odd number of plaquettes such that the FKT algorithm is not directly applicable; in that case the frustrated bond on the boundary provides an additional entropy of order $\mathcal{O}(\ln(N_{\rm bnd}))$, with $N_{\rm bnd}$ the number of boundary sites. Despite significant finite-size effects, the residual entropy clearly converges to a finite value for $N\to\infty$, which we estimate as $s_{\rm res}\equiv S_{\rm res}/N=0.102(2)$.
@@
where the number of vertices ($N$), edges ($N_{E}$) and plaquettes ($N_{P}$) of the lattice are related by Euler’s characteristic for planar graphs, $N-N_{E}+N_{P}=1$; this differs by unity from the literature since our definition of plaquettes excludes the outside. For our hyperbolic lattice we can obtain closed expressions for $N$, $N_{E}$, and $N_{P}$ based on its iterative construction. This eventually leads to [36]
@@
valid for large $n$ where $n$ denotes the number of lattice layers, and $\delta_{n}$ is one if $n$ is divisible by three and zero otherwise. The ground-state energy hence approaches (JN)(-JN) for $N\to\infty$, in agreement with the numerical results in the SM [36].
@@
We have also studied hyperbolic $\{3,7\}$ tilings with different boundary geometry. A type-B lattice is obtained by adding one site, i.e. one outer triangle, to each boundary bond of a type-A lattice. This construction generates a family of type-B tilings, Fig. 1(b). Our numerical data for type-B systems indicate that they settle into an ordered ferrimagnetic ground state, with ferromagnetic order in each layer and antiferromagnetic alignment from layer to layer, Fig. 7. The reason is energetic: Type-B systems have more boundary bonds than type-A systems, with a zigzag boundary shape, such that perfect antiferromagnetic order along the boundary coincides with ferromagnetic intralayer order, thus tipping the balance towards ordered states. More details on the behavior of type-B systems are in the SM [36].
Figure 7: NN bond correlations $\langle\sigma_{i}\sigma_{j}\rangle$ (color-coded) for $\{3,7\}$ systems at T=0.2T\!=\!0.2. (a) Type-A with $N=85$ in the spin-liquid regime. (b) Type-B with $N=50$, indicating ferrimagnetic order.
@@
Among the directions for future research are the investigation of quantum effects towards triangular hyperbolic quantum spin liquids – this suggests for instance to consider quantum dimer models on the present lattices – as well as the study of classical spin liquids in hyperbolic vector spin models. More broadly, open questions concern how lattice frustration manifests in different curved-space geometries, e.g., with black holes included [13, 42], and whether the spin liquids found here have implications for gravitational theories. We note that the finite residual entropy found here appears as an essential ingredient of the AdS/CFT correspondence: It was shown that the residual entropy of the Sachdev-Ye model [43] matches the entropy associated to the AdS2 throat arising in Reissner-Nordström charged black hole geometries [44] – a result which triggered intense research establishing holographic dualities for the Sachdev-Ye-Kitaev (SYK) model [45]. Since the Sachdev-Ye model is a spin model, our results raise the possibility that a similar duality may be realized between spin chains and antiferromagnetic Ising models. A first check here will be to compare correlation functions for spin chains at the boundary of hyperbolic spaces, as calculated e.g. in Ref. 46, with the correlation functions found here. Moreover, our results emphasize the strong influence of the boundary conditions on the ground-state degeneracy and thus on the possible presence of a gravity dual.